# Log-normal Superstatistics in the Confined Motion of Ants

### v4

@@ Abstract
We report the emergence of Log-normal Superstatistics in the collective motion of ants confined in a quasi-2D arena and exposed to a panic-inducing stimulus. A data-driven superstatistical Langevin model accurately reproduces the transition from stationary behavior to an organized escape response, characterized by non-Gaussian velocity distributions and a fluctuating diffusion coefficient. Our findings show that danger information propagates via a memory-limited, cascade-like mechanism, resulting in a stable cluster formation despite individual memory constraints. These discoveries establish a crucial connection between Superstatistics formalisms and living active matter beyond a unicellular level, and provide a foundation for the understanding of the biological origin of Log-normal type diffusion in confined environments.
Complex systems often exhibit emergent behavior that challenges classical Statistical Mechanics. In particular, far-from-equilibrium systems can feature fast local dynamics that coexist with slow fluctuations of intensive parameters such as temperature or diffusivity [1, 2, 3, 4, 5]. Superstatistics (i.e., a superposition of different local equilibrium statistics) has been successfully applied to describe these frameworks with strong parameter fluctuations, such as turbulence, convection and active particles [6, 7, 8, 9]. In living matter, devoted studies include DNA architecture and the motion of proteins and tumor cells [10, 11, 12]. Although its relevance to macroscopic biological systems has been theoretically proposed [13], it had not been directly explored experimentally. The collective behavior of animals, particularly eusocial insects like ants, presents an ideal substrate for applying superstatistical frameworks. Colonies of ants respond to environmental stimuli using distributed sensing and local communication, displaying pattern formation, foraging trails, and escape dynamics [14, 15, 16, 17].
A canonical starting point for modeling directional animal movement is the Langevin equation [18, 19, 20, 21, 13]:
@@
where $v_{i}$ is the $i$-th Cartesian component of velocity of each individual animal, $\gamma$ is a positive parameter that encloses directional persistence, and $\xi_{i}$ is a Gaussian white noise of intensity $\sigma$. Superstatistics extends this view by allowing intensive parameters such as $\beta=2\gamma/\sigma^{2}$ to fluctuate on longer time scales. The system is assumed to reach local equilibrium over a short time $\tau_{1}$, while the parameter $\beta$ varies over a slower scale $\tau_{2}\gg\tau_{1}$. Consequently, within an interval of duration $\tau_{2}$, $\beta$ can be treated as approximately constant and the velocity distribution takes the Gaussian form $p(v_{i}|\beta)=(\beta/2\pi)^{1/2}\exp{(-\beta v_{i}^{2}/2)}$ . Therefore, to capture the long-time behavior, one integrates over the ensemble of local equilibria weighted by the distribution $f(\beta)$, yielding the marginal density for the $i$-th velocity component $f(v_{i})=\int_{0}^{\infty}p(v_{i}|\beta)f(\beta)\,d\beta$ which yields diverse non-Gaussian stationary distributions documented in many physical systems [22, 23, 24, 25, 26, 27].
Here, we ask whether a fundamental statistical framework underlies ants movement in confinement and persists or adapts when ants are perturbed by panic stimuli. We show that such a framework aligns to the principles of Superstatistics and study the motion of confined ants as a case study for superstatistical analysis. Ants display non-Gaussian velocity distributions and fluctuating diffusivity, features seen in single-particle tracking of soft and biological matter [28, 29, 30, 31, 32, 33, 34]. When exposed to danger signals transmitted from a neighboring arena, ants collectively form a cluster that shifts away from the source of danger, revealing a transition from short individual memory to long-term collective behavior. We propose a super-statistical model that captures both regimes, providing what appears to be a unique experimental report of Superstatistics in the animal kingdom, and, notably, the Log-normal type in living systems.
@@
Experiments were performed in dual acrylic arenas, labelled danger and safe, of $6\times 6\times 1\,\mathrm{cm^{3}}$ connected via a Thin barrier designed to allow contacts of different types (see End Matter) (Fig. 1a). Each cell was populated with nearly 20 adult workers of the Cuban ant species Atta insularis (Fig. 6a), randomly selected from a single colony and previously studied in both natural and laboratory settings [35, 36, 37, 38, 39]. After a 5-minute baseline, $50\,\mathrm{\mu L}$ of insect repellent was introduced into the danger cell. Video data were recorded for a total of 15 minutes at 30 fps. In control (Ctrl) experiments no ants were introduced into the danger arena. Each protocol was repeated 10 times using different individuals to avoid memory effects [40].
@@
A significant shift of the center of mass (CM), measured via simple video binarization, away from the barrier was observed after introducing the repellent, indicating an escape response (Fig. 1b,d). The negligible shift in the Ctrl protocol (Fig. 1c,e) confirmed that this effect was not due to passive diffusion but driven by inter-ant communication. The CM velocity increases sharply after repellent exposure, indicating a rapid initial reaction followed by a slower phase in which ants form a stable cluster away from the barrier. This two-stage process, i.e., early activation followed by clustering, will be discussed later. Additional protocols, time-resolved dynamics, statistical validation, and experiment videos with CM evolution are provided in End Matter and Supplementary Information (SI), where we also indicate that optical signals or simple antennation are not significant mechanisms for panic transmission (Fig. 7).
@@
Figure 1: (a) Experimental setup and snapshots of experiments for (b) Thin and (c) Ctrl protocols with the center of mass (CM) trajectory colored by 5-minute intervals (green, blue, red). (d) Net horizontal CM displacement relative to its initial position for the Thin protocol, as predicted by the superstatistical model in Eq. (5) (SSModel); (e) corresponding experimental data for the Ctrl protocol.
Baseline regime was characterized by tracking 18 randomly selected ants, using the yupi library for trajectory analysis [41]. To avoid boundary effects, only trajectory segments beyond $\ell_{w}=0.6\,\mathrm{cm}$ from all walls were used in the analysis, where $\ell_{w}$ is a characteristic length obtained from the exponential decay of ant acceleration near walls (Fig. 5).
@@
Fig. 2a shows that the acceleration component parallel to the velocity averages to $\mu(v)=-\gamma v$, where $\gamma^{-1}=0.53\,\mathrm{s}$. The orthogonal component averages to zero across speeds. However, the intensity of fluctuations in both components depends nonlinearly on speed and follows the same trend within their error bars. We model these behavior with
@@
Figure 2: (a) Mean (blue/orange) and fluctuation intensity (dark/light green) of parallel and perpendicular acceleration components vs. speed. Parallel acceleration scales linearly, perpendicular remains near zero, and both speed-dependent standard deviations fit Eq. (2). Note that $\mu(v)$ and $\sigma(v)$ differ in units (i.e., $\mathrm{cm\,s}^{-2}$ and $\mathrm{cm\,s}^{-3/2}$, respectively). (b) Autocorrelation function of acceleration fluctuations around their mean (Noise Autocorrelation Function, NACF). Parallel and perpendicular components are shown in blue and orange, respectively. The fluctuations exhibit no temporal correlation, indicating uncorrelated noise. (c) Probability density function (PDF) of parallel and perpendicular acceleration fluctuations. The non-Gaussian shape is accurately captured by the SSModel (Eq. 4), which combines a Log-normal and Gaussian dynamics, but it cannot be explained only by Gaussian noise (G. Noise) with the state-dependent intensity $\sigma(v)$. (d) Time-average of local flatness (Eq. (3)). The dashed line marks the Gaussian value. Inset: typical realization of $D(t)$ (dark red) varies more slowly than $v_{x}(t)$ (dark blue) over $\sim 50\,\mathrm{s}$.
@@
The acceleration fluctuations show negligible temporal correlations (Fig. 2b). However, Fig. 2c shows that they deviate from the typical Gaussian noise of Eq. (1), and that although state-dependent noise $\sigma(v)\xi_{i}(t)$ yields non-gaussianity, it fails to explain our observations. Thus a natural step beyond constant noise intensity is to treat it as a slowly varying stochastic process, placing us in the realm of Superstatistics. Therefore, these non-Gaussian fluctuations can be explained based on a superposition of local Gaussians with variances $\beta(t)$ fluctuating over long temporal scales. In consequence, to verify time scale separation in our data, we computed the time-dependent flatness $\kappa_{i}(t)$ of velocity components [43]:
$\kappa_{i}(t)=\frac{1}{T-t}\int_{0}^{T-t}\frac{\langle(v_{i}-\langle v_{i}\rangle)^{4}\rangle}{\langle(v_{i}-\langle v_{i}\rangle)^{2}\rangle^{2}}dt^{\prime}$ (3)
@@
Here, $T$ is the total measurement time, and $\langle\cdot\rangle$ denotes a local mean over window $t$. The slow time scale $\tau_{2}$ was extracted as the window size at which the flatness $\kappa_{i}(t)$ reaches its Gaussian value, indicating local equilibrium of velocity statistics (Fig. 2d). We find $\tau_{2}=3.3\,\mathrm{s}$, which is much longer than the velocity decorrelation time $\tau_{1}=\gamma^{-1}=0.53\,\mathrm{s}$; their ratio $\tau_{2}/\tau_{1}\approx 6$ confirms the existence of an intensive parameter $\beta(t)$ that fluctuates on a timescale about 6 times slower than the velocity itself.
Figure 3: (a) Probability density (PDF) of diffusion coefficient $D$ (standardized) fitted to a Log-normal distribution. Inset shows the tail up to 15 standard deviations. (b) Velocity PDF with Gaussian, Laplace, and power-law (Lévy stable) fits. The SSModel (4) (solid line) captures the central peak and tails beyond Gaussian decay.
@@
The slowly varying stochastic process $\beta(t)$ was then computed over sliding windows of length $\tau_{2}$ as the inverse of the local velocity variance, $\beta(t)=[\langle v_{i}^{2}(t)\rangle_{\tau_{2}}-\langle v_{i}(t)\rangle_{\tau_{2}}^{2}]^{-1}$. Consequently, we explored suitable powers $p$ such that $\beta^{p}$ could make the dynamics fall into one of the universality superstatistical classes (e.g., $\chi^{2}$ or Log-normal [44]) and found that p=1p=-1 yields the diffusion coefficient $D\propto\beta^{-1}$ as the relevant slowly varying quantity. Fig. 3a shows that $D$ follows a Log-normal distribution with shape parameter $s=1.27$. Although Log-normal and stretched exponential fits can appear similar when tails are limited [44], the inset shows a robust Log-normal fit extending over 15 standard deviations. Typical time series of $v_{x}(t)$ and $D(t)$ in Fig. 2d (inset) visually confirm their separate time scales.
@@
Once gathered the measured persistence time $\gamma^{-1}$, the speed-dependent noise intensity $\sigma(v)$, and the slowly varying diffusion coefficient $D(t)$, we set them into a superstatistical Langevin model [2] for the velocity component $v_{i}(t)$ in the safe cell before panic onset:
$\frac{d}{dt}v_{i}(t)=-\gamma v_{i}(t)+\sigma(v)\sqrt{D(t)}\,\xi_{i}(t)$ (4)
@@
We refer to this as the SSModel. For simulations, at each time step, $D(t)$ and the Gaussian noise $\xi_{i}(t)$ are combined into $\eta(t)=\sqrt{D(t)}\,\xi_{i}(t)$, which is standardized and scaled by $\sigma(v)$. Numerical integration was performed using a Runge-Kutta method of a strong order 1.5 within a Milstein scheme to compute Stratonovich-type integrals [45]. Fig. 3b shows that the SSModel accurately predicts the velocity distribution: no discrepancies are observed in the shapes, consistent with the arena’s spatial homogeneity and isotropy. The model reproduces both the low-speed peak and the non-Gaussian tails through the combined effects of $\sigma(v)$ and Log-normal fluctuations in $D(t)$. It also captures the non-Gaussian fluctuations of parallel and perpendicular accelerations (Fig. 2c).

### v5

# Log-normal Superstatistics Reveals Statistical Resilience in the Panic Response of Confined Ants

@@ Abstract
We report the emergence of Log-normal Superstatistics in the collective motion of ants confined in a quasi-2D arena and exposed to a panic-inducing stimulus. A data-driven superstatistical Langevin model accurately reproduces the transition from stationary behavior to an organized escape response, characterized by non-Gaussian velocity distributions and a stochastic diffusion coefficient. Our findings show that danger information propagates via a memory-limited, cascade-like mechanism, resulting in a stable cluster formation despite individual memory constraints. These results indicate that a slowly varying diffusivity arises from the multiplicative combination of interaction-mediated processes under confinement, leading naturally to Log-normal fluctuations. The persistence of this statistical structure under panic reveals a form of collective resilience, establishing a mechanistic bridge between Superstatistics and living active matter in confined environments.
I Introduction
Complex systems often exhibit emergent behavior that challenges classical Statistical Mechanics. In particular, far-from-equilibrium systems can feature fast local dynamics that coexist with slow fluctuations of intensive parameters such as temperature or diffusivity [1, 2, 3, 4, 5]. Superstatistics (i.e., a superposition of different local equilibrium statistics) has been successfully applied to describe these frameworks with strong parameter fluctuations, such as turbulence, convection and active particles [6, 7, 8, 9]. In living matter, devoted studies include DNA architecture and the motion of proteins and tumor cells [10, 11, 12]. Although superstatistical ideas have been theoretically connected to macroscopic biological systems [13], their dedicated experimental application to animal movement is still developing [14]. The collective behavior of animals, particularly eusocial insects like ants, presents an ideal substrate for applying superstatistical frameworks. Colonies of ants respond to environmental stimuli using distributed sensing and local communication, displaying pattern formation, foraging trails, and escape dynamics [15, 16, 17, 18].
A canonical starting point for modeling directional animal movement is the Langevin equation [19, 20, 21, 22], which describes the $i$-th Cartesian velocity component $v_{i}$ of an individual as a balance between a frictional term and a stochastic noise. In its simplest form, it reads:
@@ I Introduction
where $\gamma$ is a positive parameter that encloses directional persistence, and $\xi_{i}$ is a Gaussian white noise of intensity $\sigma$. Superstatistics extends this view by allowing intensive parameters such as $\beta=2\gamma/\sigma^{2}$ to fluctuate on longer time scales. The system is assumed to reach local equilibrium over a short time $\tau_{1}$, while the parameter $\beta$ varies over a slower scale $\tau_{2}\gg\tau_{1}$. Consequently, within an interval of duration $\tau_{2}$, $\beta$ can be treated as approximately constant and the conditioned velocity distribution takes the Gaussian form:
$p(v_{i}|\beta)=\left(\frac{\beta}{2\pi}\right)^{1/2}e^{\displaystyle{-\beta v_{i}^{2}/2}}$ (2)
Therefore, to capture the long-time behavior, one integrates over the ensemble of local equilibria weighted by the distribution $f(\beta)$, yielding the marginal density for the $i$-th Cartesian velocity component:
$f(v_{i})=\int_{0}^{\infty}p(v_{i}|\beta)f(\beta)\,d\beta$ (3)
which yields diverse non-Gaussian stationary distributions documented in many physical systems [23, 24, 25, 26, 27, 28].
Here, we ask whether a fundamental statistical framework underlies ants movement in confinement and persists or adapts when ants are perturbed by panic stimuli. We show that such a framework aligns to the principles of Superstatistics and study the motion of confined ants as a case study for superstatistical analysis. To this end, we first describe the experimental setup and the emergence of the collective escape response under confinement. We then analyze a baseline regime to construct a superstatistical description of single-ant motion, and next examine how this framework extends to the panic regime and the resulting cluster dynamics. Finally, we discuss the persistence of the underlying statistical structure under perturbation and its implications for collective resilience in living active matter. The proposed superstatistical model captures both regimes, providing an experimental report of Superstatistics in the animal kingdom.
II Experimental details
@@
Experiments were performed in dual acrylic arenas, labelled danger and safe, of $6\times 6\times 1\,\mathrm{cm^{3}}$ connected via a Thin barrier designed to allow contacts of different types (see Appendix B) (Fig. 1a). Each cell was populated with nearly 20 adult workers of the Cuban ant species Atta insularis (Fig. 6a), randomly selected from a single colony and previously studied in both natural and laboratory settings [29, 30, 31, 32]. This density is consistent with crowded nest regions and dense foraging lines [33, 34], and matches occupancy regimes in which collective escape asymmetries and panic-induced effects have previously been reported under confinement [35, 36]. After a 5-minute baseline, $50\,\mathrm{\mu L}$ of insect repellent was introduced into the danger cell. Video data were recorded for a total of 15 minutes at 30 fps. In control (Ctrl) experiments no ants were introduced into the danger arena. Each protocol was repeated 10 times using different individuals to avoid memory effects [37].
@@
A significant shift of the center of mass (CM), measured via simple video binarization, away from the barrier was observed after introducing the repellent, indicating an escape response (Fig. 1b,d). The negligible shift in the Ctrl protocol (Fig. 1c,e) confirmed that this effect was not due to passive diffusion of the repellent and is consistent with interaction-mediated information transfer across the two arenas. Fig. 1d shows that the CM velocity increases sharply after repellent exposure, indicating a rapid initial reaction followed by a slower phase in which ants form a stable cluster away from the barrier. This two-stage process, i.e., early activation followed by clustering, will be discussed later. Additional protocols, time-resolved dynamics, statistical validation, and experiment videos with CM evolution are provided in Appendixes and Supplementary Information (SI), where we also indicate that optical signals or simple antennation are not significant mechanisms for panic transmission (Fig. 7).
@@
Figure 1: Experimental Setup and Collective Escape Response. (a) Experimental setup and snapshots of experiments for (b) Thin and (c) Ctrl protocols with the center of mass (CM) trajectory colored by 5-minute intervals (green, blue, red). (d) Net horizontal CM displacement relative to its initial position for the Thin protocol, as predicted by the superstatistical model in Eq. (10) (SSModel); (e) corresponding experimental data for the Ctrl protocol.
III Baseline regime
Baseline regime was characterized by tracking 18 randomly selected ants, using the yupi library for trajectory analysis [38]. The sample size was chosen to ensure reliable averages, targeting $\sim 20$ individuals as a compromise between minimal small-sample statistics ($\sim 10$) and the regime where ensemble averages begin to approach Gaussian convergence ($\sim 30$), and trajectories were selected by uniform random sampling among ants visible in the safe cell at the beginning of the baseline window across the 10 repetitions. To avoid boundary effects, only trajectory segments beyond $\ell_{w}=0.6\,\mathrm{cm}$ from all walls were used in the analysis, where $\ell_{w}$ is a characteristic length obtained from the exponential decay of ant acceleration near walls (Fig. 5).
Figure 2: Speed-Dependent Fluctuations and Time-Scale Separation in the Baseline Regime. (a) Mean (blue/orange) and fluctuation intensity (dark/light green) of parallel and perpendicular acceleration components vs. speed. Error bars account for the standard error of the mean. Parallel acceleration scales linearly, perpendicular remains near zero, and both speed-dependent standard deviations fit Eq. (4). Note that $\mu(v)$ and $\sigma(v)$ differ in units (i.e., $\mathrm{cm\,s}^{-2}$ and $\mathrm{cm\,s}^{-3/2}$, respectively). (b) Autocorrelation function of acceleration fluctuations around their mean (Noise Autocorrelation Function, NACF). Parallel and perpendicular components are shown in blue and orange, respectively. The fluctuations exhibit no temporal correlation, indicating uncorrelated noise. (c) Probability density function (PDF) of parallel and perpendicular acceleration fluctuations. The non-Gaussian shape is accurately captured by the SSModel (Eq. 8), which combines a Log-normal and Gaussian dynamics, but it cannot be explained only by Gaussian noise (G. Noise) with the state-dependent intensity $\sigma(v)$. (d) Time-average of local flatness of Cartesian velocity components (Eq. (5)). The dashed line marks the Gaussian value. Inset: typical realization of $D(t)$ (dark red) varies more slowly than $v_{x}(t)$ (dark blue) over $\sim 50\,\mathrm{s}$.
@@ III Baseline regime
Fig. 2a shows that the acceleration component parallel to the velocity averages to $\mu(v)=-\gamma v$, where $\gamma^{-1}=0.53\,\mathrm{s}$. The perpendicular component averages to zero across speeds. However, the intensity of fluctuations in both components depends nonlinearly on speed and follows the same trend within their error bars. We model these behavior with
@@
The acceleration fluctuations show negligible temporal correlations (Fig. 2b). However, Fig. 2c shows that they deviate from the Gaussian noise assumed in Eq. (1). Although state-dependent noise $\sigma(v)\,\xi_{i}(t)$ introduces non-Gaussian features, it does not fully account for the observed tails. A natural extension beyond time-invariant noise intensity is therefore to treat it as a slowly varying stochastic process, placing the system within the realm of Superstatistics. In this picture, the observed non-Gaussian fluctuations arise from a superposition of local Gaussian processes whose variances fluctuate over longer timescales.
To test whether such a separation of timescales exists in the data, we computed the time-dependent flatness $\kappa_{i}(t)$ of the Cartesian velocity components [40]:
$\kappa_{i}(t)=\frac{1}{T-t}\int_{0}^{T-t}\frac{\langle(v_{i}-\langle v_{i}\rangle)^{4}\rangle_{t}}{\langle(v_{i}-\langle v_{i}\rangle)^{2}\rangle_{t}^{2}}\,dt^{\prime},$ (5)
where $T$ is the total measurement time, $\langle\cdot\rangle_{t}$ denotes averages computed within sliding windows of duration $t$, and $t^{\prime}$ labels the starting time of the window. The flatness is the normalized fourth moment (kurtosis) and equals 3 for a Gaussian distribution; deviations from this value quantify non-Gaussianity. Here, the window size $t$ plays the role of a coarse-graining timescale: for small $t$, statistics reflect instantaneous fluctuations, while increasing $t$ progressively averages over slower modulations.
@@
The slow time scale $\tau_{2}$ was extracted as the window size at which the flatness $\kappa_{i}(t)$ approaches the Gaussian value, indicating that, despite the presence of multiplicative noise, at this coarse-graining scale the local velocity statistics are well approximated by a quasi-Gaussian distribution (Fig. 2d). We find $\tau_{2}=3.3\,\mathrm{s}$, which is much longer than the velocity decorrelation time $\tau_{1}=\gamma^{-1}=0.53\,\mathrm{s}$; their ratio $\tau_{2}/\tau_{1}\approx 6$ confirms the existence of an intensive parameter $\beta(t)$ that fluctuates on a timescale about 6 times slower than the velocity itself.
The slowly varying stochastic process $\beta(t)$ was then computed over sliding windows of length $\tau_{2}$ as the inverse of the local velocity variance,
$\beta(t)=[\langle v_{i}^{2}(t)\rangle_{\tau_{2}}-\langle v_{i}(t)\rangle_{\tau_{2}}^{2}]^{-1}$ (6)
@@
Consequently, we explored suitable powers $p$ such that $\beta^{p}$ could make the dynamics fall into one of the universality superstatistical classes (e.g., $\chi^{2}$ or Log-normal [41]) and found that p=1p=-1 yields the diffusion coefficient $D\propto\beta^{-1}$ as the relevant slowly varying quantity. Fig. 3a shows that $D$ follows a Log-normal distribution with PDF
$f(D)=\frac{1}{\sqrt{2\pi s^{2}}\,D}\exp\!\left(-\frac{\ln^{2}D}{2s^{2}}\right)$ (7)
where the fitted shape parameter is $s=1.27$. Although Log-normal and stretched exponential fits can appear similar when tails are limited [41], the inset shows a robust Log-normal fit extending over 15 standard deviations. Typical time series of $v_{x}(t)$ and $D(t)$ in Fig. 2d (inset) visually confirm their separate time scales.
Figure 3: Log-normal Diffusivity and Non-Gaussian Velocities in the Baseline Regime. (a) Probability density (PDF) of diffusion coefficient $D$ (standardized) fitted to a Log-normal distribution. Inset shows the tail up to 15 standard deviations. (b) Velocity PDF with Gaussian, Laplace, and power-law (Lévy stable) fits. The SSModel (8) (solid line) captures the central peak and tails beyond Gaussian decay.
@@
Once gathered the measured persistence time $\gamma^{-1}$, the speed-dependent noise intensity $\sigma(v)$, and the slowly varying diffusion coefficient $D(t)$, we set them into a superstatistical Langevin model [2] for a Cartesian velocity component $v_{i}(t)$ in the safe cell before panic onset:
$\frac{d}{dt}v_{i}(t)=-\gamma v_{i}(t)+\sigma(v)\sqrt{\hat{D}(t)}\,\xi_{i}(t)$ (8)
where $\hat{D}(t)$ is introduced through the normalization of the compound noise generated by $D(t)$. More explicitly, what we use from $D(t)$ in the simulations is its Log-normal shape parameter $s$, which determines the statistics of the non-Gaussian noise $\eta(t)=\sqrt{D(t)}\,\xi_{i}(t)$. This compound noise is then standardized to unit variance, yielding $\hat{\eta}(t)=\sqrt{\hat{D}(t)}\,\xi_{i}(t)$, and finally scaled by $\sigma(v)$. We refer to this as the SSModel.
@@
Numerical integration was performed using a Runge-Kutta method of a strong order 1.5 within a Milstein scheme to compute Stratonovich-type integrals [42]. Fig. 3b shows that the SSModel accurately predicts the velocity distribution: no discrepancies are observed in the shapes, consistent with the arena’s spatial homogeneity and isotropy. The model reproduces both the low-speed peak and the non-Gaussian tails through the combined effects of $\sigma(v)$ and Log-normal fluctuations in $D(t)$. It also captures the non-Gaussian fluctuations of parallel and perpendicular accelerations (Fig. 2c).
IV Panic regime
Panic regime was examined by tracking 48 trajectories of ants during the next 10 minutes following repellent injection. A larger sample ($\sim 50$ trajectories) was targeted to ensure sufficient statistics for time-resolved estimates, and manual tracking yielded 48 usable trajectories.